# Quantifying Resolution Limits in Pedestal Profile Measurements with Gaussian Process Regression

### v1

@@ 3 GPR as Low-Pass Filtering
In this section, we show how GPR can be interpreted as a form of low-pass filtering, and use this to give an explicit procedure for interpreting and choosing hyperparameters in the Matérn family of kernels, including the squared exponential kernel as a limit.
@@ 3.1 Spectral Analysis of GPR Inference
$H(\xi)$ is also called the equivalent kernel [19].
@@
This behavior for $H(\xi)$ indicates that GPR inference $\hat{f}(x)$ provides an estimate for a low-pass filtered version of $f$, not $f$ itself. Low frequency components of $f$ are inferred in an approximately unbiased way, but high frequency components of $f$ are attenuated by the bias. This phenomenon can be interpreted as a simple form of spectral bias, observed in other machine learning settings [35].
@@
where $C_{\nu}$ is the same as in (6). $S$ is the signal-to-noise rate, which is the signal-to-noise ratio $\sigma^{2}/\sigma_{\varepsilon}^{2}$, multiplied by the sampling rate normalized to the length scale of the kernel $\ell/\Delta x$. Plots of the filter cutoff as a function of the signal-to-noise rate for different values of $\nu$ are shown in figure 2. We remark that $H(\xi)\propto\xi^{-(2\nu+1)}$ for $\xi\gg\xi_{*}$, suggesting that $12\nu+$ 1 can be interpreted as the order of the low-pass filter.
@@ 3.3 Recommendations from Low-Pass Interpretation
Figure 3: Illustration of ‘ringing’ artifacts caused by the Gibbs phenomenon when using GPR to infer non-smooth functions. True profile (blue line) and data (blue scatter) are shown along with inferences (orange with ±$\pm 34\%$ confidence interval shaded) overplotted. The inverse filter cutoff $1/\xi_{*}$ is also shown.
One key observation from the preceding analysis is that the characteristic length scale of the inferred profile is not simply given by $\ell$, and instead is determined by the cutoff frequency $\xi_{*}$ which depends both on $\ell$ and the other hyperparameters through the signal-to-noise rate $S$. Intuitively, since GPR doesn’t constrain the functional form of the resulting fit, functions with high-frequency components above the length scale $\ell$ can be inferred if the quality of the data is sufficiently high.
@@
    Smoothness $\nu$: If the inferred function lacks discontinuities in any of its derivatives, $\nu=\infty$ should be chosen since it has the sharpest drop-off in filtering high-frequency components. However, if there is expected to be an approximate discontinuity of any of the derivatives, such as due to the presence of a localized transport barrier, finite $\nu$ can be considered. For problems with discontinuous derivative, $\nu=2.5$ was found to strike a good balance between alleviating the Gibbs phenomenon while retaining a reasonable drop-off rate for high frequency noise.
@@
    Length scale $\ell$: For a given variance $\sigma^{2}$ and smoothness $\nu$, the cutoff frequency $\xi_{*}$ can be computed in terms of $\ell$ using (7). Aliasing usually should be avoided, so the cutoff frequency must satisfy $\xi_{*}\Delta x\lesssim 1/2$. The exact choice of $\ell$ can then be chosen based on the resolution requirements through $\xi_{*}$.
In practice, real experimental data often is given on irregularly spaced locations and with non-identical error bars. In this case, it is still possible to compute a signal-to-noise rate $\bar{S}=\sigma^{2}\ell^{2}/\overline{\sigma_{\varepsilon}^{2}}\overline{\Delta x}$ by using the average spacing $\overline{\Delta x}$ and average measurement variance $\overline{\sigma_{\varepsilon}^{2}}$ for the data points in the region of interest. We show later in section 5 how using this average signal-to-noise rate allows the idealized low-pass analysis to be used to inform analyses in real-world scenarios.
@@ 4 Information Theory for GPR
While in the previous section we developed a theory for GPR in an idealized case of filtering continuous functions, in this section we focus on information theory to answer the question: without making any simplifying assumptions, exactly how many measurements contribute to the inference of a value at a given point? We begin by introducing the concept of Fisher information and its relation to inference problems in section 4.1. This is followed by the definition of $N_{eff}$, a metric that measures the effective number of measurements which contribute to an inference, in section 4.2. Finally, we discuss some mathematical properties and examples of $N_{eff}$ in section 4.3, and we discuss the relationship between $N_{eff}$ and robustness of GPR inference to unmodeled errors in section 4.4.
@@ 4.1 Fisher Information
where $p(\mathbf{y};\boldsymbol{\theta})$ is the PDF of $\mathbf{Y}$ for a given value of the parameter $\boldsymbol{\theta}$ [36]. Note the Fisher information is a positive semi-definite matrix.
@@
To interpret second key property, note that the eigenvalues of the covariance matrix give a measure of the spread of the distribution of $\mathbf{T}(\mathbf{Y})$, with smaller eigenvalues implying that the distribution of $\mathbf{T}(\mathbf{Y})$ is more closely concentrated around the mean $\boldsymbol{\theta}$. One consequence of the matrix inequality above is that the smallest eigenvalue of $\operatorname{Cov}[\mathbf{T}(\mathbf{Y})]$ is bounded below by the inverse of the largest eigenvalue of $\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$, putting a limit on how tightly concentrated the distribution can be.
@@
The so-called data processing inequality [37] says that $\mathcal{I}_{\mathbf{T}(\mathbf{Y})}[\boldsymbol{\theta}]\leq\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$, which reduces this to the earlier expression. When $\mathcal{I}_{\mathbf{T}(\mathbf{Y})}[\boldsymbol{\theta}]=\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$ this tells us that our estimator is using all available Fisher information from $\mathbf{Y}$ to inform our knowledge of $\boldsymbol{\theta}$. Note that the Fisher information, and more generally information geometry theory, has seen usage in plasma physics in both statistical modeling and inference [38, 39].
@@ 4.2 Definition of $N_{eff}$
¯$\bar{\mathcal{I}}_{\mathbf{Y}}[\bar{f}(x)]:=\sum_{i=1}^{N}\mathcal{I}_{Y_{i}}[\bar{f}(x)]\left(\tfrac{\mathcal{I}_{Y_{i}}[\bar{f}(x)]}{\mathcal{I}_{\mathbf{Y}}[\bar{f}(x)]}\right)$ (12)
@@
¯$\bar{\mathcal{I}}_{\mathbf{P}}[\bar{f}(x)]=\sum_{j=1}^{M}\mathcal{I}_{P_{j}}[\bar{f}(x)]\left(\tfrac{\mathcal{I}_{P_{j}}[\bar{f}(x)]}{\mathcal{I}_{\mathbf{Y}}[\bar{f}(x)]}\right)$
@@ 4.3 Mathematical Properties and Simple Example of $N_{eff}$
$N_{\mathbf{Y}}[\bar{f}(x)]=\left(\frac{\left\lVert\mathbf{s}\right\rVert_{1}}{\left\lVert\mathbf{s}\right\rVert_{2}}\right)^{2}$
@@
    (Boundedness) $1\leq N_{eff}\leq N$. $N_{eff}=1$ if and only if exactly one $s_{i}$ is non-zero, and $N_{eff}=N$ if and only if all of the $s_{i}$ are non-zero and equal.
@@
Property 1 shows that $N_{eff}$ respects the physical number of measurements, and cannot be below 1 or above $N$. Property 2 shows when $R$ measurements contribute equally to an inference, $N_{eff}=R$. Property 3 is somewhat more subtle, and shows that $N_{eff}$ gives a metric which is independent of the size of the error bars. Roughly speaking, the Fisher informtion scales like the inverse of the variance, a result which is exact for Gaussian random variables. The scaling invariance says that $N_{eff}$ is insensitive to the absolute magnitude of the information, and is only sensitive to the relative contribution of the different measurements. An inference with small error bars may still lack credibility if it only depends on a small number of measurements, as the magnitude of the error bars of the underlying data may be underestimated due to errors not captured in $\varepsilon_{i}$.
@@ 4.4 $N_{eff}$ Controls Unmodeled Errors
We close with an important application of $N_{eff}$ for controlling unmodeled errors, which are uncertainties in the data not captured by the noise covariance $\Sigma_{\varepsilon}$. Often, not all sources of error are either explicitly or correctly accounted for in the model of the measurement noise $\varepsilon_{i}$. A common example of practical interest are systematic uncertainties in the measurements due to calibration errors. In this section, we demonstrate how $N_{eff}$ gives a rigorous bound to the degree to which these unmodeled errors can propagate into GPR inferences, with larger $N_{eff}$ guaranteeing that the unmodeled error will make a smaller impact.
To make this statement quantitative, suppose the measured data $\tilde{Y}_{i}$ differ from the idealized measurements $Y_{i}$ by an unmodeled error $\delta_{i}$, so $\tilde{Y}_{i}=Y_{i}+\sigma_{i}\delta_{i}$. For simplicity we assume the measurement errors $\varepsilon_{i}$ are independent of both each other and the unmodeled errors with variance $\sigma_{i}^{2}$. The unmodeled errors are thus normalized to $\sigma_{i}$.
Using $\tilde{\mathbf{Y}}$ in place of $\mathbf{Y}$ in (2) would produce a profile $\tilde{f}(x)$ which differs from the original inference $\hat{f}(x)$ by
$\tilde{f}(x)-\hat{f}(x)=\sum_{i=1}^{N}\beta_{i}(x)\sigma_{i}\delta_{i}$
@@
$\tilde{\Delta}(x):=\frac{\operatorname{Var}[\tilde{f}(x)-\hat{f}(x)]}{\bar{\sigma}^{2}_{\varepsilon}(x)}$ (15)
@@
$\tilde{\Delta}(x)$ then quantifies the mean square deviation of $\tilde{f}(x)$ from $\hat{f}(x)$ over different realizations of the unmodeled noise $\delta_{i}$, assuming that $\mathbb{E}[\delta_{i}]=0$. Note that similar definitions can be given for the derivative by replacing $\beta_{i}(x)$ by $\beta_{i}^{\prime}(x)$.
Now, suppose the unmodeled errors $\delta_{i}$ are statistically independent. Then,
$\tilde{\Delta}(x)^{2}\leq\frac{||\boldsymbol{\delta}||_{4}^{4}}{N_{\mathbf{Y}}[f(x)]}$ (16)
That is, the square of the normalized deviation is necessarily smaller than $\ell^{4}$ norm of the vector of normalized errors $\boldsymbol{\delta}$, divided by $N_{eff}$. Thus, if the magnitude of the unmodeled error is fixed by $||\boldsymbol{\delta}||_{4}$, then one way to decrease the resulting impact on the inference is to increase $N_{eff}$.
A proof of the inequality (16), which is a straightforward application of the Cauchy-Schwarz inequality for matrix inner products, is given in C. In the case where measurements in a given bin have correlated unmodeled error, the binned version $N_{\mathbf{P}}$ of $N_{eff}$ can be used in place of $N_{\mathbf{Y}}$.

### v2

@@ 3 GPR as Low-Pass Filtering
In this section, we show how GPR can be interpreted as a form of low-pass filtering, and use this to give an explicit procedure for interpreting and choosing ranges of hyperparameters in the Matérn family of kernels, including the squared exponential kernel as a limit.
@@ 3.1 Spectral Analysis of GPR Inference
$H(\xi)$ is also called the equivalent kernel [rasmussen_gaussian_2008].
@@
This behavior for $H(\xi)$ indicates that GPR inference $\hat{f}(x)$ provides an estimate for a low-pass filtered version of $f$, not $f$ itself. Low frequency components of $f$ are inferred in an approximately unbiased way, but high frequency components of $f$ are attenuated by the bias. This phenomenon can be interpreted as a simple form of spectral bias, observed in other machine learning settings [rahaman_spectral_2019]. An alternative interpretation of these results is that $\hat{f}(x)$ is an unbiased estimator for the low-pass filtered version of $f$.
@@
where $C_{\nu}$ is the same as in (7). $S$ is the signal-to-noise rate, which is the signal-to-noise ratio $\sigma^{2}/\sigma_{\varepsilon}^{2}$, multiplied by the sampling rate normalized to the length scale of the kernel $\ell/\Delta x$. Plots of the filter cutoff as a function of the signal-to-noise rate for different values of $\nu$ are shown in figure 2. We remark that $H(\xi)\propto\xi^{-(2\nu+1)}$ for $\xi\gg\xi_{*}$, suggesting that $2\nu+1$ can be interpreted as the order of the low-pass filter. We also note that a completely flat pass-band with exact cutoff would be achievable with a sinc kernel [tobar_band-limited_2019], which may be useful in some applications.
@@ 3.3 Recommendations from Low-Pass Interpretation
Figure 3: Illustration of ‘ringing’ artifacts caused by the Gibbs phenomenon when using GPR to infer non-smooth functions. True profile (blue line) and data (blue scatter) are shown along with inferences (orange with $\pm 34\%$ confidence interval shaded) overplotted. The inverse filter cutoff $1/\xi_{*}$ is also shown.
One key observation from the preceding analysis is that the characteristic length scale of the inferred profile is not simply given by $\ell$, and instead is determined by the cutoff frequency $\xi_{*}$ which depends both on $\ell$ and the other hyperparameters through the signal-to-noise rate $S$. Intuitively, since GPR doesn’t constrain the prior functional form of the inference, functions with high-frequency components above the length scale $\ell$ can be inferred if the quality of the data is sufficiently high.
@@
    Smoothness $\nu$: If the inferred function lacks sharp changes in any of its derivatives, $\nu=\infty$ should be chosen since it has the sharpest drop-off in filtering high-frequency components. However, if a sharp change in any of the derivatives is expected, such as due to the presence of a localized transport barrier, finite $\nu$ can be considered. For these problems, $\nu=2.5$ was found to strike a good balance between alleviating the Gibbs phenomenon while retaining a reasonable drop-off rate for high frequency noise.
@@
    Length scale $\ell$: For a given variance $\sigma^{2}$ and smoothness $\nu$, the cutoff frequency $\xi_{*}$ can be computed in terms of $\ell$ using (8). Aliasing should usually be avoided, so the cutoff frequency must satisfy $\xi_{*}\Delta x\lesssim 1/2$. The exact choice of $\ell$ can then be chosen based on the resolution requirements through $\xi_{*}$.
To avoid over-regularization for pedestal profiles, it is reasonable to assert that profile features on the scale of the pedestal itself need be inferred. This corresponds to a requirement that $\xi_{*}w_{ped}\gtrsim 1$ for some measure $w_{ped}$ of pedestal width. The estimate of the pedestal width may itself depend on the degree of smoothing applied, so satisfying this requirement may involve a degree of exploration of the hyperparameter space, which we consider later in section 5.
Figure 4: Illustration of effective low-pass filter behavior of GPR when applied to non-uniformly sampled data with non-identical error bars. Figure labels, as well as hyperparameters chosen for the GPR kernel, are the same as in Figure 3.
In practice, real experimental data is often given on irregularly spaced locations and with non-identical error bars. One way to apply the low-pass filter analysis is to compute an average signal-to-noise rate
$\bar{S}=\sigma^{2}\ell^{2}/\overline{\sigma_{\varepsilon}^{2}}\overline{\Delta x}$ (10)
using the average spacing $\overline{\Delta x}$ and average measurement variance $\overline{\sigma_{\varepsilon}^{2}}$ for the data points in the region of interest. A caveat for this approach is that the low-pass filter analogy and corresponding filter cutoff $\xi_{*}$ will only give an approximation of the actual behavior of the GPR fit. Figure 2 and equation (8) show that the scaling of $\xi_{*}$ with $S$ is either a fairly weak power law or logarithmic. For $S\gg 1$, which is the typical case, the dependence of $\xi_{*}$ on $\nu$ is much more impactful, so (8) may still give the correct overall trend of $\xi_{*}$ using $\bar{S}$, even in the case with irregularly spaced locations and non-identical error bars.
In figure 4, we demonstrate an example of a GPR inference which randomizes the measurement locations and error bars of the synthetic data from the example in figure 3, while keeping the average signal-to-noise ratio fixed. Again, note that no noise has been applied to the synthetic data to emphasize the effect of the low-pass filter. The inferences do not deviate significantly from the regularly sampled identical measurement case shown in figure 3, showing the relevance of the low-pass filter analogy when using an averaged signal-to-noise rate.
To close this section, we briefly discuss how this picture in terms of resolution requirements relates to the Bayesian approach involving hyperpriors. The design of hyperpriors typically requires the incorporation of physical assumptions or intuition built across a particular class of profiles. For a given hyperprior, evaluating the cutoff frequency across the range of posterior hyperparmeters gives a quantitative picture of how assumptions on the frequency content of $f$ will bias the resulting estimate $\hat{f}$.
If the computed cutoff frequencies do not meet the resolution requirements, one of three approaches can be taken to address issues of over-regularization depending on the circumstances. The first is to attempt increase the signal-to-noise rate, in which case the formula for the cutoff frequency can give an order of magnitude estimate for the additional amount of data necessary to meet the resolution requirement. If the set of measurements if fixed, the second approach is to see if the hyperprior can be redesigned to incorporate less restrictive hypotheses on the expected frequency content of the profiles, in which case the formula for the cutoff frequency gives a guide on how changes in the hyperparmeters will affect the resulting inference. If neither approach works to address over-regularization, then the final approach is to conclude that the measurements cannot resolve the full desired range of frequencies. Recalling that GPR gives an unbiased estimate for the low-pass filtered profile, the inferred profile can be presented along with the cutoff frequency as an estimate of the degree of over-regularization applied by the inference process.
@@ 4 Information Theory for GPR
While in the previous section we developed a theory for GPR in an idealized case of filtering continuous functions, in this section we focus on information theory to answer the question: without making any simplifying assumptions, exactly how many measurements contribute to the inference of a value at a given point?
We begin by introducing the concept of Fisher information and its relation to inference problems in section 4.1. This is followed by the definition of $N_{eff}$, a metric that measures the effective number of measurements which contribute to an inference, in section 4.2. Finally, we discuss some mathematical properties and examples of $N_{eff}$ in section 4.3, and we discuss the relationship between $N_{eff}$ and robustness of GPR inference to unmodeled errors in section 4.4.
@@ 4.1 Fisher Information
where $p(\mathbf{y};\boldsymbol{\theta})$ is the PDF of $\mathbf{Y}$ for a given value of the parameter $\boldsymbol{\theta}$ [lehmann_theory_1998]. Note the Fisher information is a positive semi-definite matrix.
@@
To interpret the second key property, note that the eigenvalues of the covariance matrix give a measure of the spread of the distribution of $\mathbf{T}(\mathbf{Y})$, with smaller eigenvalues implying that the distribution of $\mathbf{T}(\mathbf{Y})$ is more closely concentrated around the mean $\boldsymbol{\theta}$. One consequence of the matrix inequality above is that the smallest eigenvalue of $\operatorname{Cov}[\mathbf{T}(\mathbf{Y})]$ is bounded below by the inverse of the largest eigenvalue of $\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$, putting a limit on how tightly concentrated the distribution can be.
@@
The so-called data processing inequality [zamir_proof_1998] says that $\mathcal{I}_{\mathbf{T}(\mathbf{Y})}[\boldsymbol{\theta}]\leq\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$, which reduces this to the earlier expression. When $\mathcal{I}_{\mathbf{T}(\mathbf{Y})}[\boldsymbol{\theta}]=\mathcal{I}_{\mathbf{Y}}[\boldsymbol{\theta}]$ this tells us that our estimator is using all available Fisher information from $\mathbf{Y}$ to inform our knowledge of $\boldsymbol{\theta}$. Note that the Fisher information, and more generally information geometry theory, has seen usage in plasma physics in both statistical modeling and inference [kim_time-dependent_2020, kim_probabilistic_2025].
@@ 4.2 Definition of $N_{eff}$
$\bar{\mathcal{I}}_{\mathbf{Y}}[\bar{f}(x)]:=\sum_{i=1}^{N}\mathcal{I}_{Y_{i}}[\bar{f}(x)]\left(\tfrac{\mathcal{I}_{Y_{i}}[\bar{f}(x)]}{\mathcal{I}_{\mathbf{Y}}[\bar{f}(x)]}\right)$ (14)
@@
$\bar{\mathcal{I}}_{\mathbf{P}}[\bar{f}(x)]=\sum_{j=1}^{M}\mathcal{I}_{P_{j}}[\bar{f}(x)]\left(\tfrac{\mathcal{I}_{P_{j}}[\bar{f}(x)]}{\mathcal{I}_{\mathbf{Y}}[\bar{f}(x)]}\right)$
@@ 4.3 Mathematical Properties and Simple Example of $N_{eff}$
$N_{\mathbf{Y}}[\bar{f}(x)]=\left(\frac{\left\lVert\mathbf{s}\right\rVert_{1}}{\left\lVert\mathbf{s}\right\rVert_{2}}\right)^{2}$ (17)
@@
    (Boundedness) If at least one $s_{i}$ is non-zero, $1\leq N_{eff}\leq N$. Furthermore, $N_{eff}=1$ if and only if exactly one $s_{i}$ is non-zero, and $N_{eff}=N$ if and only if all of the $s_{i}$ are non-zero and equal.
@@
Property 1 shows that $N_{eff}$ respects the physical number of measurements, and cannot be below 1 or above $N$. Property 2 shows when $R$ measurements contribute equally to an inference, $N_{eff}=R$. Property 3 is somewhat more subtle, and shows that $N_{eff}$ gives a metric which is independent of the size of the error bars. Roughly speaking, the Fisher information scales like the inverse of the variance, a result which is exact for Gaussian random variables. The scaling invariance says that $N_{eff}$ is insensitive to the absolute magnitude of the information, and is only sensitive to the relative contribution of the different measurements. An inference with small error bars may still lack credibility if it only depends on a small number of measurements, as the magnitude of the error bars of the underlying data may be underestimated due to errors not captured in $\varepsilon_{i}$.
We remark that property 3 also shows that $N_{eff}$ is not a proxy for the total amount of information contributed by the measurements. If that is desired, the total information $\mathcal{I}_{\mathbf{Y}}[\bar{f}(x)]$ should be used directly instead. The choice to normalize the total information by the weighted average information $\bar{\mathcal{I}}_{\mathbf{Y}}[\bar{f}(x)]$ at the same location $x$ leads to the scaling invariance, but other choices may lead to other metrics capturing other information-theoretic properties of the inference.
@@ 4.4 $N_{eff}$ Controls Unmodeled Errors
We close with an important interpretation of $N_{eff}$ involving the control of unmodeled errors, which are uncertainties in the data not captured by the noise covariance $\Sigma_{\varepsilon}$. Often, not all sources of error are explicitly or correctly accounted for in the model for the measurement noise $\varepsilon_{i}$. A common example of practical interest are systematic uncertainties in the measurements due to calibration errors. In this section, we demonstrate how $N_{eff}$ gives a rigorous bound to the degree to which a particular class unmodeled errors can propagate into GPR inferences, with larger $N_{eff}$ guaranteeing that the unmodeled error will make a smaller impact.
We consider a simple model of the impact of unmodeled correlated error between different measurements. As before, let $\mathbf{P}=P_{1},...,P_{M}$ be a partition of $\{1,...,N\}$. Suppose $\boldsymbol{\delta}=[\delta_{1},...,\delta_{M}]^{T}$ is a random vector of deviations. We assume that each of the deviations $\delta_{j}$ are independent from the other deviations and from the measurement errors $\boldsymbol{\varepsilon}$. We also assume the deviations have zero mean and finite variance, although we do not require the deviations to be normally distributed.
For each bin $P_{j}$, we suppose that each measurement $\tilde{Y}_{i},i\in P_{j}$ differs from the idealized measurement $Y_{i}$ by
$\tilde{Y}_{i}=Y_{i}+\sigma_{i}\delta_{j}$
In other words, for each bin $P_{j}$, the measurement error for $i\in P_{j}$ is equal to $\tilde{\varepsilon}_{i}=\varepsilon_{i}+\sigma_{i}\delta_{j}$. Using $\tilde{\mathbf{Y}}$ in place of $\mathbf{Y}$ in (2) would produce a profile $\tilde{f}(x)$ which differs from the idealized inference $\hat{f}(x)$ by
$\tilde{f}(x)-\hat{f}(x)=\sum_{j=1}^{M}\left[\delta_{j}\sum_{i\in P_{j}}\beta_{i}(x)\sigma_{i}\right]$
@@
$\tilde{\Delta}(x):=\frac{\operatorname{Var}[\tilde{f}(x)-\hat{f}(x)|f]}{\bar{\sigma}^{2}_{\varepsilon}(x)}$ (18)
@@
$\tilde{\Delta}$ quantifies the mean squared deviation between the inferred profile $\tilde{f}$ with an idealized inference $\hat{f}$ which has the deviation $\boldsymbol{\delta}$ subtracted out. From a frequentist view, we can imagine this variance as computed over an ensemble of experiments which measure an identical plasma profile $f$, where each experiment has a different realization of the noise $\boldsymbol{\varepsilon}$ and systematic deviation $\boldsymbol{\delta}$. Note that similar definitions can be given for the derivative by replacing $\beta_{i}(x)$ by $\beta_{i}^{\prime}(x)$.
Then, this mean squared deviation satisfies
$\tilde{\Delta}(x)^{2}\leq\frac{\overline{D^{2}}}{N_{\mathbf{P}}[\bar{f}(x)]}$ (19)
where $\overline{D^{2}}:=\sum_{j=1}^{M}|P_{j}|^{2}\operatorname{Var}[\delta_{j}]^{2}$ is the sum of the squared variances weighted by the number of elements $|P_{j}|$ in each bin. That is, given a fixed magnitude of the deviations as quantified by $\overline{D^{2}}$, the resulting deviation in the inference $\tilde{f}-\hat{f}$ can be reduced by increasing $N_{eff}$. Note that this result does not hold for systematic deviations which are correlated between different bins; in that case, a different binning must be used.
The proof of (19) is a straightforward application of the Cauchy-Schwarz inequality, which we show in C. We also discuss an analogous result for the derivative, and give a simple example that achieves the worst case predicted by the bound in the appendix. We remark that the origin of the squared variance scaling comes from the choice to weight the average information per measurement by the information itself which leads to $\bar{\mathcal{I}}_{\mathbf{P}}\sim\mathcal{I}_{P_{j}}^{2}$. Different choices of weighting could potentially lead to other information-theoretic inequalities which control the unmodeled error. Other choices for capturing the unmodeled error may also suggest different information-theoretic metrics to consider. We leave systematic study of these choices and their empirical behavior to future work.